Assume there was a situation in which you observed 50 of these binomial random variables, and they were all mutually independent. That is, assume i... Xso Binomial(n,p). Find an approzimate distribution for R = Σ the Write the name of the distribution and identify its parameter(s). Justify all claims and steps
Answer #1

Result:- We\ will\ use\ the\ result\ of CLT\ here.\ If, X_1,X_2,....X_m\ are\ n\ iid\ random\newline observations,\ such\ that\ E[X]=\mu,V[X]=\sigma^2,Then,\frac{\sqrt m(\bar X-\mu)}{\sigma}\sim N(0,1)\newline such\ that,\bar X\sim N(\mu,\frac{\sigma^2}{m})\newline Here,E[X_i]=np,V[X_i]=np(1-p),m=50,So,\bar{X}\sim N(np,\frac{np(1-p)}{50}).\newline So, parameters\ of\ Normal\ distribution\ are,\mu=np,\sigma^2=\frac{np(1-p)}{50}